OpenStudy (lani):

hi, i am trying to fin the derivative of y= pi^2 + e^x + x^2 + x^x^(1/2) at x=1. Help please!! :)

OpenStudy (anonymous):

Which part do you not know how to differentiate?

OpenStudy (amistre64):

\[y= \pi^2 + e^x + x^2 + x^{x^(1/2)}\] this?

OpenStudy (amistre64):

pi^2 is a constant it goes to 0 e^x derives to e^x x^2 derives to 2x and x^(x^(1/2)) is the tricky one

OpenStudy (amistre64):

\[x^{x^{1/2}}=x^{(1/2)x}\]

OpenStudy (anonymous):

\[y=\pi^2+e^x+x^2+x^\sqrt{x}\] First, remove the constants, then individually differentiate each part separated by + signs. Differentiate each part individually, and use logarithmic differentiation for the last part.

OpenStudy (amistre64):

id assume it goes: \[Dx(x^{(1/2)x}) \implies \frac{1}{2}*\frac{x}{2}(x)^{\frac{x-2}{2}}\]

OpenStudy (amistre64):

\[\frac{x*x^{(x-2)/2}}{4}\] and the top simplifies further since like bases add exponents

OpenStudy (amistre64):

but if x = 1; we dont need to go thru all the window dressing do we :)

OpenStudy (anonymous):

\[y=x^\sqrt{x}\] \[\ln(y)=\sqrt{x}\ln(x)\] \[Dx(\ln(y)=\sqrt{x}\ln(x))\] \[y'/y = 1/2 *1/\sqrt{x} * \ln(x) + \sqrt{x}/x \] \[y' = (1/2 *1/\sqrt{x} * \ln(x) + \sqrt{x}/x) /y \] \[y' = (1/2 *1/\sqrt{x} * \ln(x) + \sqrt{x}/x) /x^\sqrt{x} \] Into which, you could plug 1.